OpenStudy (anonymous):

please help medals given find the integral of log(1+sinx) from -pi/2 to pi/2

OpenStudy (anonymous):

\[\int\limits_{-\pi/2}^{\pi/2}\ln (1+sinx)dx\]

OpenStudy (anonymous):

what method are you supposed to use? you will not find an anti derivative for this one

OpenStudy (anonymous):

i do not know how to solve it i suppose using integration by parts would and the answer is given as (-pi)*(ln2) but i dont know how to reach there

OpenStudy (anonymous):

must be some trick, because there is no closed form anti derivative for this one

OpenStudy (anonymous):

damn, i didn't even notice it is an improper integral too

OpenStudy (anonymous):

ya maybe i grew the graph for this function maybe that could help